OpenStudy (anonymous):

Frank kicks a soccer ball off the ground and in the air with an initial velocity of 30 feet per second. Using the formula H(t) = -16t^2 + vt + s, what is the maximum height the soccer ball reaches? 14.1 feet 13.7 feet 13.2 feet 15.2 feet

OpenStudy (anonymous):

1 don't know what 1'm doing wrong 1 keep getting 0.

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

What is the "v" and "s" in the formula for H(t)?

OpenStudy (anonymous):

velocity and speed 1 think.

OpenStudy (asnaseer):

Is this all the information given to you in the question?

OpenStudy (anonymous):

yes. You're supposed to use the quadratic formula

OpenStudy (asnaseer):

If "v" is supposed to be the initial velocity then you can re-write this as:\[H(t)=-16t^2+30t+s\]That still leaves the unknown "s" in this equation. Usually "s" is used to denote distance, but here you are give H(t) to represent the "distance" in terms of height. Is this question from a text book?

OpenStudy (anonymous):

no the question is on my homework.

OpenStudy (anonymous):

s would be 0

OpenStudy (asnaseer):

I am getting a lot of "lag" on this site right now - so sorry if I seem to be taking a long time to answer. Why do you say s=0?

OpenStudy (anonymous):

that's just how my teacher told me to set it up. s represents how far from the ground the starting point is.

OpenStudy (asnaseer):

Ok, so we actually have:\[H(t)=-16t^2+30t\]correct?

OpenStudy (anonymous):

yes but the 0 is necessary for the quadratic formula

OpenStudy (asnaseer):

Why do you think you need to use the quadratic formula? The quadratic formula is usually used to solve equations like:\[15t^2+12t-3=0\]Where as here you are being asked to find the maximum value of H given:\[H(t)=-16t^2+30t\]Have you studied calculus yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this is algebra 1 honors. 1'm in 8th grade.

OpenStudy (asnaseer):

Have you been taught about how to rewrite a quadratic in vertex form?

OpenStudy (anonymous):

1'm not very good at it.

OpenStudy (anonymous):

a(x - h)^2 + k right?

OpenStudy (asnaseer):

Ok - but it seems that you have been taught that method. I therefore think that you should use this method to solve this. Yes - that is the vertex form

OpenStudy (anonymous):

\[h=\frac{ -b }{ 2a }\] \[k= c- \frac{ 2b }{ 4a }\] ?

OpenStudy (asnaseer):

If you look at the equation given to you, we have:\[H(t)=-16t^2+30t\]So first step would be to factor out the 16 to give:\[H(t)=16\left(-t^2+\frac{30t}{16}\right)\]

OpenStudy (asnaseer):

we can now factor the terms inside the braces to get something like \((x-h)^2\) can you see how to do that?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

It's called completing the square

OpenStudy (anonymous):

okay

OpenStudy (asnaseer):

what do you get?

OpenStudy (anonymous):

1 have no idea. sorry. 1'm horrible at math. : /

OpenStudy (asnaseer):

np - lets take a look at an example to help you...

OpenStudy (asnaseer):

lets say we have:\[x^2+\color{red}{8}x\]the first thing we would do is look at the coefficient of \(x\) and halve it. In this case we would halve \(\color{red}{8}\) to get \(\color{red}{4}\). so we then calculate what this equals:\[(x+\color{red}{4})^2\]

OpenStudy (anonymous):

oh okay, 1 see

OpenStudy (asnaseer):

sorry - my internet went down

OpenStudy (anonymous):

1ts fine

OpenStudy (asnaseer):

ok - if we now expand this we get:\[(x+\color{red}{4})^2=x^2+\color{red}{8}x+\color{blue}{16}\]so you can see if get an extra term of \(\color{blue}{16}\) we therefore need to subtract this in order to get our original equation, i.e.:\[x^2+\color{red}{8}x=(x+\color{red}{4})^2-\color{blue}{16}\]

OpenStudy (asnaseer):

If we now go back to your equation, we have:\[H(t)=16\left(-t^2+\frac{30t}{16}\right)\]and we need to complete the square on:\[-t^2+\frac{30t}{16}\]

OpenStudy (asnaseer):

here we have to be careful because we have a negative sign on the \(t^2\) term

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

so first we can rewrite this as:\[-\left(t^2-\frac{30t}{16}\right)\]so we need to complete the square on:\[t^2-\frac{30t}{16}\]understand so far?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so, we you recall my explanation above, we need to halve the coefficient of \(t\) in this case and write:\[(t-\frac{15}{16})^2\]

OpenStudy (asnaseer):

because \(\frac{15}{16}\) is half of \(\frac{30}{16}\)

OpenStudy (anonymous):

okay

OpenStudy (asnaseer):

if we now expand this we get:\[(t-\frac{15}{16})^2=t^2-\frac{30t}{16}+\color{blue}{\left(\frac{15}{16}\right)^2}\]to get our desired terms we need to subtract the blue term. therefore:\[t^2-\frac{30t}{16}=(t-\frac{15}{16})^2-\color{blue}{\left(\frac{15}{16}\right)^2}\]

OpenStudy (asnaseer):

making sense?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good. now we can go back to our original equation which was:\[H(t)=16\left(-t^2+\frac{30t}{16}\right)=16\left(-\left(t^2-\frac{30t}{16}\right)\right)\]\[=16\left(-\left((t-\frac{15}{16})^2-\color{blue}{\left(\frac{15}{16}\right)^2}\right)\right)\]\[=16\left(-(t-\frac{15}{16})^2+\color{blue}{\left(\frac{15}{16}\right)^2}\right)\]

OpenStudy (asnaseer):

If you look at this equation can you see what the maximum value would be?

OpenStudy (asnaseer):

we can rewrite this slightly better as:\[H(t)=16\left(\color{blue}{\left(\frac{15}{16}\right)^2}-\color{red}{\left(t-\frac{15}{16}\right)^2}\right)\]

OpenStudy (asnaseer):

since the terms inside the outer braces are both squared, the maximum must be when the red term is zero.

OpenStudy (asnaseer):

since the blue term is a constant and the red term varies with \(t\)

OpenStudy (anonymous):

okay

OpenStudy (asnaseer):

so we get the maximum height when \(t=\frac{15}{16}\) (which would make the red term equal to zero) and the value of the maximum would be:\[H_{max}=16\left(\color{blue}{\left(\frac{15}{16}\right)^2}\right)=\cancel{16}\times\frac{15\times15}{\cancel{16}\times16}=\frac{225}{16}\approx14.1\text{ feet}\]

OpenStudy (anonymous):

geez, it seems like a lot of work for an answer.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (asnaseer):

I took extra time to ensure you understood each step

OpenStudy (asnaseer):

once you get good at completing the square then you should be able to work through these types of problems very quickly

OpenStudy (asnaseer):

yw :)